# PaECTER: Patent-level Representation Learning using Citation-informed Transformers

### v1

@@ 5 Conclusion
The PaECTER model can be a useful tool for researchers working with patent text. For example, researchers often rely on explicit citations to determine relatedness or knowledge flows. Not only are citations rare, but they can also be selected in a strategic or opportunistic manner (Jaffe and de Rassenfosse,, 2017).
In response to the limitations of citation-based analysis, researchers increasingly turn to textual analysis to determine relatedness. A notable instance in this evolving field is the work of (Kelly et al.,, 2021), who identified breakthrough patents by assessing their textual dissimilarity to the existing patent corpus, employing a TF-IDF transformation of patent texts. While informative, this method has computational constraints.
The PaECTER model addresses these challenges by enabling the establishment of patent relatedness in scenarios where citations are absent or may be strategically manipulated. Its application offers immediate and scalable benefits, making it a powerful tool for exploring patent landscapes and discerning innovative breakthroughs. This capability significantly contributes to the nuanced understanding of knowledge flows and patent ecosystem dynamics, thus broadening the horizons for patent research and analysis.
The models are publicly available on Huggingface:
6 Acknowledgements
Our work has been enabled by grant 2023/8401 from the European Patent Office under their Academic Research Programme (EPO-ARP) 2021. We also thank the team at the Max Planck Computing and Data Facility (MPCDF) for allowing us to use their high-performance computing cluster.

### v2

@@ 5 Conclusion
The PaECTER model provides a valuable resource for researchers and practitioners analyzing patent text. This includes applications that go beyond traditional TF-IDF transformations or aim to identify similar patents during the examination process.
PaECTER enables the ranking of patent documents by similarity independent of citations. This capability contributes to a deeper understanding of knowledge flows and patent system dynamics, and offers promising directions for future research and analysis in the patent domain. The training and test datasets can also be used for advancing research and development in patent analytics.
The model and training data are publicly available on Huggingface:
6 Limitations
The PaECTER model has several limitations that are important to consider:
Token length constraints.
The underlying BERT architecture processes sequences of up to 512 tokens. As a result, PaECTER was trained on patent titles and abstracts, rather than full descriptions or claims, which are central to defining a patent’s legal scope. While the model captures general semantic aspects of patents, it may be less effective at representing their fine-grained technical details.
Language scope.
The model was trained exclusively on English-language patent texts. It is therefore not directly applicable to non-English patents. Users working with non-English texts should apply high-quality machine translation prior to inference.
Temporal relevance.
Because the training data reflects existing and past technologies, the model may underperform on emerging innovations not yet well-represented in the dataset. In particular, it may struggle to identify or classify technologies requiring new Cooperative Patent Classification (CPC) categories in the future. This is left for future analysis.
Acknowledgements
Our work has been enabled by grant 2023/8401 from the European Patent Office under their Academic Research Programme (EPO-ARP) 2021. We also thank the team at the Max Planck Computing and Data Facility (MPCDF) for allowing us to use their high-performance computing cluster. Finally we thank participants of the European Patent Office’s ARP Workshop 2023 (in particular our discussant, Gaétan de Rassenfosse) as well as the 2024 Munich Summer Institute for helpful comments.